# General effective reduction theory of integral polynomials of given non-zero discriminant and its applications

### v1

@@ D. Distribution of monogenic number fields
As is well-known, all quadratic number fields are monogenic. For degree $n=3$, the first example of a non-monogenic number field was given by Dedekind (1878). For certain values $n\geq 3$, there are various results of the shape that there are infinitely many non-monogenic number fields of degree $n$.
@@
Akhtari (2020) showed that a positive proportion of cubic number fields, when ordered by their absolute discriminant, are not monogenic. Alpöge, Bhargava and Shnidman (2020) proved more precisely that, if isomorphism classes of cubic number fields are ordered by their absolute discriminant, then a positive proportion are not monogenic and yet have no local obstruction to being monogenic. Recently, Alpöge, Bhargava and Shnidman (2024) proved a similar result for quartic number fields, and even that a positive proportion of quartic number fields are not rationally monogenic.
@@
For $n=3,4,6$, tables of Gaál (2019) show that the frequency of monogenic number fields $K$ of degree $n$ is decreasing in tendency as the absolute value of the discriminant $|D_{K}|$ increases.
@@
Denote by $N_{n}(X)$ the number of isomorphism classes of monogenic number fields $K$ of degree $n$ with $|D_{K}|\leq X$ and with associated Galois group $S_{n}$.
Theorem D.1 (Bhargava, Shankar and Wang, 2022).
For every $n\geq 2$ we have
$N_{n}(X)\gg X^{{1/2+1}/(n-1)}\ \ \text{as }X\to\infty.$
In the same paper, the authors conjecture that $X^{{1/2+1}/(n-1)}$ is the right order of magnitude.
E. Arithmetic characterization of monogenic and multiply monogenic number fields
@@
In this direction there are many important results for deciding the monogenicity or non-monogenicity of number fields from certain special infinite classes, including cyclotomic, abelian, cyclic, pure, composite number fields, certain quartic, sextic, multiquadratic number fields and relative extensions, and parametric families of number fields defined by binomial, trinomial,… irreducible polynomials.
In their proofs various types of tools are used, among others Dedekind’s criterion; Newton polygons; Montes’ algorithm; Ore’s theorem; Engström’s theorem; Gröbner basis approach; reduction to binomial Thue equations; irreducible monic polynomials with square-free discriminant; non-squarefree discriminant approach; infinite parametric families of number fields; use of the index form approach.
For details, we refer to Dedekind (1878) and to the books Hensel (1908), Hasse (1963), Narkiewicz (1974), Evertse and Győry (2017), Gaál (2019) and the references given there. For some recent developments, see also the survey article Gaál (2024) with many interesting special results. We note that Hasse’s problem is not yet solved in full generality.
In general, while the monogenicity can be, its multiplicity cannot be determined by the arithmetic methods utilized.
Problem 5.
Give an arithmetic characterization of multiply monogenic number fields.
Clearly, Hasse’s problem and Problem 5 do not properly belong to the reduction theory of integral polynomials.
Dedekind’s necessary criterion for monogenicity of a number field was generalized by Del Corso, Dvornicich and Simon (2005) to a criterion for rational monogenicity. Perhaps this provides a tool to construct more examples of number fields that are not rationally monogenic.
Acknowledgment. We are very grateful to Dr. Csanád Bertok for typing a substantial part of the manuscript.

### v2

@@ B. Distribution of monogenic and non-monogenic number fields
As is well-known, all quadratic number fields and cyclotomic fields are monogenic. For degree $n=3$, the first example of a non-monogenic number field was given by Dedekind (1878). For every $n\geq 3$, there are infinitely many isomorphism classes of monogenic, cf. Kedlaya (2012), and infinitely many isomorphism classes of non-monogenic number fields of degree $n$.
Let $K$ be a number field, and $\{1,\omega_{2},\ldots,\omega_{n}\}$ a $\mathbb{Z}$-module basis of $\mathcal{O}_{K}$. Denote by $I(X_{2},\ldots,X_{n})$ the associated index form, as introduced in Subsection 5.4. Thus, if $\alpha=x_{1}+x_{2}\omega_{2}+\cdots+x_{n}\omega_{n}$ with $x_{1},\ldots,x_{n}\in\mathbb{Z}$, then $\mathcal{O}_{K}:\mathbb{Z}[\alpha]]=|I(x_{2},\ldots,x_{n})|$. Consequently, $K$ is monogenic if and only if $I(x_{2},\ldots,x_{n})=\pm 1$ is solvable in $x_{2},\ldots,x_{n}\in\mathbb{Z}$. We say that $K$ has no local obstruction to being monogenic if for every prime number $p$, the equation $I(x_{2},\ldots,x_{n})=\pm 1$ has a solution $x_{2},\ldots,x_{n}$ in the $p$-adic integers. This notion does not depend on the choice of $\omega_{2},\ldots,\omega_{n}$. We recall some recent results.
Theorem B.1 (Alpöge , Bhargava and Shnidman, 2025).
@@
Let $K$ run through the isomorphism classes of cubic fields, ordered by their absolute discriminant. Then a positive proportion of them are not monogenic, and yet have no local obstruction to being monogenic.
Subsequently, but published earlier, the same authors proved the following result for quartic fields. Recall that a number field is called rationally monogenic if its ring of integers is rationally monogenic.
Theorem B.2 (Alpöge, Bhargava and Shnidman, 2024).
Let $K$ run through the isomorphism classes of quartic fields, ordered by their absolute discriminant. Then a positive proportion of them are not rationally monogenic, and yet have no local obstruction to being monogenic.
@@
For $n=3,4,6$, tables of Gaál (2019) suggest that the density of monogenic number fields $K$ of degree $n$ decreases with the absolute value of the discriminant $|D_{K}|$.
Bhargava, Shankar and Wang established the following pioneering result.
Theorem B.3 (Bhargava, Shankar and Wang, 2022).
@@
Denote by $M_{n}(X)$ the number of isomorphism classes of monogenic number fields $K$ of degree $n$ with $|D_{K}|\leq X$ and with associated Galois group $S_{n}$. Then for every $n\geq 2$ we have
$M_{n}(X)\gg X^{1/2+1/n}\ \ \text{as }X\to\infty.$
The authors conjecture that the exponent on $X$ is optimal.
In Part II of their paper, the authors proved a corresponding result for rationally monogenic number fields:
Theorem B.4 (Bhargava, Shankar and Wang, 2025).
Denote by $RM_{n}(X)$ the number of isomorphism classes of rationally monogenic number fields $K$ of degree $n$ with $|D_{K}|\leq X$ and with associated Galois group $S_{n}$. Then for every $n\geq 3$ we have
$RM_{n}(X)\gg X^{1/2+1/(n-1)}\ \ \text{as }X\to\infty.$
Let $N_{n}(X)$ denote the number of isomorphism classes of number fields $X$ of degree $n$ with $|D_{K}|\leq X$. It is conjectured that $N_{n}(X)\asymp X$ as $X\to\infty$. This is easy for $n=2$. Davenport and Heilbronn (1971) proved this conjecture for $n=3$ and Bhargava (2005, 2010) for $n=4,5$.
C. Arithmetic characterization of monogenic and multiply monogenic number fields
@@
In this direction there are many important results for deciding the monogenicity or non-monogenicity of number fields from certain special infinite classes, including quadratic, cyclotomic, abelian, cyclic, pure, composite number fields, certain quartic, sextic, multiquadratic number fields and relative extensions, and parametric families of number fields defined by binomial, trinomial,… irreducible polynomials.
In their proofs various types of tools are used, among others Dedekind’s criterion; Newton polygons; Montes’ algorithm; Ore’s theorem; Engström’s theorem; Gröbner basis approach; reduction to binomial Thue equations; elliptic curve approaches, irreducible monic polynomials with square-free discriminant; non-squarefree discriminant approach; infinite parametric families of number fields; use of the index form equation approach with “small" solutions.
For details, we refer to Dedekind (1878) and to the books Hensel (1908), Hasse (1963), Narkiewicz (1974), Evertse and Győry (2017), Gaál (2019) and the references given there. For some recent developments, see also the survey article Gaál (2024) with many interesting special results, and the recent interesting papers Kaur, Kumar and Remete (2025), Sharma and Sarma (2025), Guàrdia and Perdet (2025), Gaál (2025), Harrington and Jones (2025), Yakkou, Aghzer and Boua (202?), and König (2025). We note that Hasse’s problem has not yet been solved in full generality.
A more precise version of Hasse’s problem is as follows.
Problem 7.
For $m\geq 1$, give an arithmetic characterization of those number fields whose ring of integers is $m$ times monogenic.
Clearly, Hasse’s problem and Problem 7 do not properly belong to the reduction theory of integral polynomials.
Dedekind’s necessary condition for monogenicity of a number field was generalized by Del Corso, Dvornicich and Simon (2005) to a condition for rational monogenicity. Perhaps this provides a tool to construct more examples of number fields that are not rationally monogenic.